OpenStudy (baldymcgee6):

Find the value of this limit:

OpenStudy (anonymous):

tell me the question please

OpenStudy (baldymcgee6):

\[\lim_{x \rightarrow 8} \frac{ (x^\frac{ 2 }{ 3 } - 4) }{ (x^\frac{ 1 }{ 3 } - 2) }\]

OpenStudy (anonymous):

can u attach it with clear pls

OpenStudy (baldymcgee6):

attach it with what? @sudharsa

OpenStudy (baldymcgee6):

http://gyazo.com/9e646627ecf766a91ef117db60d3c64d

OpenStudy (across):

Hint: factor the numerator.

OpenStudy (baldymcgee6):

how do we factor it? what do we take out?

OpenStudy (baldymcgee6):

@across ?

OpenStudy (across):

The numerator has the form \(a^2-b^2=(a+b)(a-b)\).

OpenStudy (anonymous):

dvide numerator and denominator by x^1/3 yeah

OpenStudy (across):

You don't have to divide, @sudharsa.

OpenStudy (anonymous):

ok ur also goog idea that mke it simple

OpenStudy (anonymous):

across is good then answer is simple yeah

OpenStudy (baldymcgee6):

but what is sqrt(x^2/3)

OpenStudy (anonymous):

if x= 8 then cubic root of it will be 2 then proceed yeah

OpenStudy (baldymcgee6):

@across I don't understand how to factor the numberator..

OpenStudy (across):

If you let \(u=x^{1/3}\), then your numerator will become \(u^2-4\). Can you factor that?

OpenStudy (baldymcgee6):

its x^(2/3), but nonetheless, (u+2)(u-2)

OpenStudy (across):

You have to let \(u=x^{1/3}\) because then \(u^2=x^{2/3}\)... now re-substitute and solve your exercise.

OpenStudy (baldymcgee6):

okay thank you.